Solve charpit method: 2z+p^2+qy+2y^2=0
1
Expert's answer
2020-04-29T18:30:02-0400

Let "f(x,y,z,p,q) \equiv 2z+p^2+qy+2y^2=0" ............(1)


Charpit's auxiliary equation are



"\frac{dp}{f_x+pf_z}=\frac{dq}{f_y+qf_z}=\frac{dz}{-pf_p-qf_q}=\frac{dx}{-f_p}=\frac{dy}{-f_q}"


....................(2)

From (1), "f_x=0, \ f_y=q+4y, \ f_z=2, \ f_p=2p,\ f_q=y" .......(3)

Using (3),(2) reduce to



"\frac{dp}{2p}=\frac{dq}{3q+4y}=\frac{dz}{-2p^2-qy}=\frac{dx}{-2p}=\frac{dy}{-y}"


Taking 1st and 4th fraction we get,


"\implies \frac{dp}{2p}=\frac{dx}{-2p}"



So that , "P=-x" ..............(4)

Now taking 2nd and 5th fraction ,we get


"\frac{dq}{3q+4y}=\frac{dy}{-y}"




"\implies \frac{dq}{dy}=-\frac{3q+4y}{y}"



"\implies \frac{dq}{dy}+\frac{3}{y}q=-4"


Which is a linear equation in first order.


"\therefore I.F=e^{\int \frac{3}{y} dy } =e^{3logy}=e^{log {y^3} }=y^3"


Now , "q \ y^3=\int (-4)y^3dy =-y^4"

"\implies q=-y" .


Now putting the value of p and q in

"dz=pdx+qdy" we get "dz=-xdx-ydy"

"\implies 2z=-x^2-y^2+c"

Where "c" is a integration constant.